. Elasticity (related to Total Revenue): (Please show your steps & decision rules)

The demand schedule for a specific computer software is as follows

Price (per copy of software)

Quantity demanded (in million)

525

68

550

64

590

62

612

60

680

55

750

48

900

43

935

40

1010

37

  1. At what price is total revenue a maximum and what quantity sold? Show your calculation/numbers
  2. When the price falls from $750 to $590, is the demand for this software elastic or inelastic or others? Explain why.
  3. When the price increases from $680 to $935, is the demand for this software elastic or inelastic or others? Explain why.
Answer #1

A) TR= P*Q

It is maximum when P = 680

Price QD TR
525 68 35700
550 64 35200
590 62 36580
612 60 36720
680 55 37400
750 48 36000
900 43 38700
935 40 37400
1010 37 37370

B) E = [Q2-Q1/(Q2+Q1/2)]/[P2-P1/(P2+P1/2)]

= [62-48/(62+48/2)]/[590-750/(590+750/2)]

= [14/55]/[160/670]

= 0.2545/0.2388

= 1.06

The demand is elastic as the value is more than one.

C) E= [40-55/(40+55/2)] / [935-680/(935+680/2)]

= [15/47.5] / [255/807.5]

= 0.3158/0.3158

= 1

The demand is unitary elastic as the value is equal to 1.